OpenStudy (anonymous):

A rod has length 3 meters and radius (1/sqrt(pi)) meters. At a distance x meters from its left end, the density of the rod is given by delta(x)= 4+5x g/m^3. Find the total mass

OpenStudy (anonymous):

We will begin by dividing the rod into into many infinitesimal segments of length dx. The mass of each segment is dm

OpenStudy (anonymous):

u understand?

OpenStudy (anonymous):

Basically I'm trying to figure out whether this is a one dimensional, 2 d or 3d

OpenStudy (anonymous):

the density varies with only x, so this is essentially 1-d

OpenStudy (anonymous):

Well it's the radius portion of it that's throwing me off

OpenStudy (anonymous):

Do you get how we divide the rod into small discs

OpenStudy (anonymous):

yeah i do. I find integral from 0 to 3 for 4+5x dx

OpenStudy (anonymous):

but that's not the right answer

OpenStudy (anonymous):

Recall that 4+5x is the DENSITY. dm = Density x length x Area

OpenStudy (anonymous):

integral from 0 to 3 of (4+5x)(pi)(r^2)dx

OpenStudy (anonymous):

yeah i know but r = 1/sqrt(pi) so (1/sqrt(pi))^2 = 1/pi so 1/pi * pi = 1

OpenStudy (anonymous):

which leaves us with integral from 0 to 3 for 4+5x dx

OpenStudy (anonymous):

34.5

OpenStudy (anonymous):

yeah that's what i got, but that's not the right answer

OpenStudy (anonymous):

whts the right ans?

OpenStudy (anonymous):

i don't know haha, that's why i'm seeking help. I've been working on this one for a while and I can't figure it out

OpenStudy (anonymous):

Well...what you did seems legit to me

OpenStudy (anonymous):

yeah me too. At least that's what I thought

OpenStudy (anonymous):

what would be the units? I found 34.5 g

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

haha never mind it worked this time. I guess it didn't like 69/2. Stupid program